Az első valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Angyal Dóra  −  −  +  −  −  +
Bakó Ágnes  −  ±  −  −  −  +
Bathó Gábor  −  +  ±  −  −  −
Bócsi Judit  +  +  +  −  +  +
Boda Judit  +  +  +  ±  +  +
Ficzere Éva  +  +  −  −  +  +
Fista Róbert  −  +  +  −  ±  +
Gyöngyösi Levente  −  −  ±  −  +  −
Helmeczi József  −  −  −  −  −  −
Jónás Ágnes  ±  +  +  −  ±  +
Kalóz Margit  +  +  −  −  +  +
Karácsony Szilveszter  −  −  ±  −  −  +
Kis Mirella  −  −  +  −  −  −
Kovács Viktor  −  −  −  −  −  −
Ködöböcz József  −  −  −  −  −  −
Lengyel Sándor  +  +  ±  −  ±  +
Mályi Nikoletta  +  +  +  +  −  +
Miskolczi Panna  +  +  ±  −  ±  +
Nagy Beáta  −  −  +  −  ±  +
Nagy Eszter  −  −  −  −  −  +
Papp Ágnes  ±  −  −  −  −  +
Pósa Csenge  −  −  −  −  +  −
Szeifert Diana  ±  −  +  −  −  −
Szpodnyi Marianna  +  +  −  −  −  +
Tóth Edit  ±  −  +  −  +  +
Tóth Veronika  +  ±  −  −  −  +